OpenStudy (anonymous):

Can someone please help me with this problem? I am BEYOND confused! :/ A 1.8 tall basketball player attempts a goal 11 from the basket that is 3.05 high. f he shoots the ball at a 47◦ angle, at what initial speed must he throw the basketball so that it goes through the hoop without striking the backboard? The acceleration of gravity is 9.81 m/s2 . Answer in units of m/s. Btw this is a vector problem!

OpenStudy (anonymous):

This is what the actual problem looks like- picture included.

OpenStudy (anonymous):

well i solved this question today :) its a 2D motion so we need to find the equation of X and Y so i got the answer as 11m/s but always consider a calc mistake so i might be wrong :)

OpenStudy (anonymous):

how did you do it?

OpenStudy (anonymous):

=..= im to lazy to think noW .. sowwiie >_ <

OpenStudy (anonymous):

is it projectial question

OpenStudy (anonymous):

yes @uzumakhi

OpenStudy (anonymous):

This equation requires 1 kinematic equation, the ones that deal with displacement. Since this is a 2-dimensional problem, I will separate the equations into their x and y components, which will leave a total of 2 equations, which I will label and use to help you solve your question. 1. \[x _{t} = x _{o} + v _{o _{x}}t\] 2. \[y _{t} = y _{o} + v _{o _{y}}t + \frac{ 1 }{ 2 }a _{y}t\] For the x direction, there is no change in acceleration, it only has initial velocity, so therefore we can ignore the acceleration part of the equation. For the y direction, there is acceleration due to gravity, so we have to keep those there. Now because this is a 2 dimensional problem, we will have to separate the x and y components for the initial velocity. To do so is simple, just requires a bit of trigonometry. Since the angle is 47 degrees, you can assume that alpha in all future equations will be that. The y component for the initial velocity is as follows:\[v _{o} \cos \left( \alpha \right) = v_{o_{x}}\] The x component for the initial velocity is as follows:\[v _{o} \sin \left( \alpha \right) = v_{o_{y}}\] You can plug these in for the initial velocity for both the x and y equations, then substitute time t with the x equation, by solving for t in equation 1. \[\frac{ x _{t} }{ v _{o _{x}} } = t\] Since equation 2 also has a value for t, you can plug it in, which should give you \[y _{t} = y _{o} + v _{o _{y}}\left( \frac{ x _{t} }{ v _{o _{x}} } \right) + \frac{ 1 }{ 2 }a _{y}\left( \frac{ x _{t} }{ v _{o _{x}} } \right)\] Now all you have to do is solve for the initial velocity, v sub zero.

OpenStudy (anonymous):

Wow that took forever to make hahaha

OpenStudy (anonymous):

o_0

OpenStudy (anonymous):

The answer that @AskSamphysics provided is roughly correct, just due to calculator error just as he stated, and rounding. But that is the same answer I got as well.

OpenStudy (anonymous):

okay! @hkim THANK YOU SO MUCH! You're an angel!

OpenStudy (anonymous):

X}

OpenStudy (anonymous):

Not a problem @natasha.aries Just give me a Best Response vote and close the question if I answered it to your needs =]

OpenStudy (anonymous):

im going to work it out again and make sure i fix my mistake! and then ill close it:)

OpenStudy (anonymous):

(\_/) (O_O) .(v v)

OpenStudy (anonymous):

No problem. Just let me know if you have any other questions while solving this problem.

OpenStudy (anonymous):

will do! :)

OpenStudy (anonymous):

Oh I just noticed a quick mistake, for the two trigonometric velocity equations, the first one is for the y component, and the second one is for the x component, though the equations themselves have those flipped. Just make sure you do the algebra right, and you should have no problem.

OpenStudy (anonymous):

im a little confused about the equations bc when i plug it in it confuses me :/

OpenStudy (anonymous):

Which one confuses you? The cosine alpha one is for the y component and the sine alpha one is for the x component

OpenStudy (anonymous):

i dont understand how i would solve for it bc i know the degrees but thats it

OpenStudy (anonymous):

Basically, you would just replace all the y components of the initial velocity with the trigonometric equation for that component. You are only solving for the following\[v _{o}\] and that is it. I guess I'll write out another question so you can see it a bit easier.\[y _{t} = \left( v _{o}\sin \left( \alpha \right) \right) t - \frac{ 1 }{ 2 } g t ^{2}\] For all values of t, you just replace it with the following. \[t=\frac{ x _{t} }{ v _{o}\cos \left( \alpha \right) } \] I'll give you a hint that should help you as well. \[\frac{ v _{o}\sin \left( \alpha \right) }{ v _{o}\cos \left( \alpha \right) } = \tan \left( \alpha \right) \]

OpenStudy (anonymous):

thank youu

OpenStudy (anonymous):

Let me know if you still need help solving the problem, this is the last one I'll make sure is solved before I retire for the night =]

OpenStudy (anonymous):

lol no go to sleep! if anything ill send it and you can answer it tomorrow!

OpenStudy (anonymous):

No worries, this was my plan anyways. I'm learning physics as well on my own, so this is good learning experience for me as well.

OpenStudy (anonymous):

are you sure? idk how long it will take! i would much rather you sleep first! and what grade are you actually in then?

OpenStudy (anonymous):

I am a college dropout, studying physics for both my own benefit and my future. I wish to be a physics professor someday =P. Therefore, no need to worry about when I'm going to sleep.

OpenStudy (anonymous):

i hope your wish comes true! :)

OpenStudy (anonymous):

Well I have to take it a step at a time. And the step I'm taking right now is helping you solve this problem, so please do not hesitate to ask if you need further assistance.

OpenStudy (anonymous):

Okay!

OpenStudy (anonymous):

what is xt in xt/vox=t ?

OpenStudy (anonymous):

That is the x position that the ball has traveled in time t. According to the question, that is basically how far the ball traveled along the x-axis, until it reaches the goal.

OpenStudy (anonymous):

how do i find that?

OpenStudy (anonymous):

It is given to you in the problem. Look closely, you should be able to see it, how far the ball traveled in the x-direction.

OpenStudy (anonymous):

11

OpenStudy (anonymous):

Correct! How about the y-direction? Make sure to take into account the height of the player as well.

OpenStudy (anonymous):

what would i put for vo?

OpenStudy (anonymous):

That is what you are finding out for. Vo is the answer to the question, and you will need algebra to get to it.

OpenStudy (anonymous):

then what is t?

OpenStudy (anonymous):

t is irrelevant. That is the time that it takes for the ball to reach there, but with the given values, is unable to be solved for. Luckily, if you solve one of the first 2 equations for t, then plug that in for t, then you won't ever need to use it.

OpenStudy (anonymous):

okay i think im close! whats yo? 0?

OpenStudy (anonymous):

Correct! Same for Xo. With the way that we setup the problem, the initial position of both the x and y coordinates of the ball is 0. You can say the ball starts at the origin.

OpenStudy (anonymous):

i got 90. 9 :/

OpenStudy (anonymous):

Most likely, the algebra was done incorrectly. What values did you get for all the variables?

OpenStudy (anonymous):

(.73135)(16.1291)+1/2 (9.81)(16.1291)

OpenStudy (anonymous):

Ah, I see where the issue was. You have to realize that since the direction that the Y axis is increasing is up, then that means gravity, would have to be, in fact, negative, for the answer to come up correctly. Try putting a - in front of the g(9.81) to solve for the problem.

OpenStudy (anonymous):

i tried that afterwards and got -67.32

OpenStudy (anonymous):

I think you might have done the algebra wrong. \[y=x \tan \left( \alpha \right) + \frac{ 1 }{ 2 }\frac{ g x ^{2} }{\left( v _{o}\cos \left( \alpha \right) \right)^{2}}\] is where you start with. From there, if you do the algebra right, you should get as follows \[\frac{ 2\left( y-x \tan \left(\alpha \right) \right) }{ g x ^{2} } = \frac{ 1 }{ \left( v _{o} \cos \left( \alpha \right) \right) ^{2}}\] if you take inverse square root, you get \[v _{o} = \frac{ \sqrt{\frac{ g x ^{2} }{ 2\left( y - x \tan \left( \alpha \right) \right) }} }{ \cos \left( \alpha \right) } \]Just solve for Vo in that equation. Plug in everything as you have it. If you get a negative value, you most likely put the wrong sign on gravity.